How many pounds of an alloy containing 60% copper must be melted with an alloy containing 20% copper to obtain 4 pounds of an alloy containing 35% copper?
Answer by checkley77(12844)

.60x+.20(4-x)=.35*4
.60x+.8-.20x=1.4
.40x=1.4-.8
.40x=.6
x=.6/.4
x=1.5 pounds of 60% copper is needed.
4-1.5=2.5 pounds of 20% copper is needed.
Proof:
.60*1.5+.20*2.5=.35*4
.9+.5=1.4
1.4=1.4